OpenStudy (anonymous):

find the perimeter of a triangle with vertices (2,4), (-1,3) and (2,-3)

OpenStudy (rogue):

Use the distance formula to find the distance between each coordinates and add them up.\[d = \sqrt{ (x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

\[\sqrt{{(-1-2)^{2}+(3-4)^{2}}}=\sqrt{9+1}=\sqrt{10}\] \[\sqrt{(2+1)^{2}+(-3-3)^{2}}=\sqrt{9+36}=\sqrt{45}=3\sqrt{5}\] \[\sqrt{(2-2)^{2}+(-3-4)^{2}}=\sqrt{49}=7\] These are the three sides, AB, BC, and AC. Now just add all 3 of them: \[\sqrt{10}+3\sqrt{5}+7\] units.

OpenStudy (anonymous):

thanks! how do i show my work on graph paper?

OpenStudy (rogue):

Plot each of the points and measure the distance between them. Or if you want, plot each of the points and use the Pythagoeran theorem to find the distance between them, which is pretty much using the distance formula.

OpenStudy (anonymous):

thank you so much! thats very helpful :)